# Counterfactual Explanations for Integer Optimization Problems

### v1

@@ 4.1 Weak Counterfactual Explanations
If $\hat{a}$ and $\hat{b}$ are fixed, the mutable parameter space is denoted by $\mathcal{H}_{c}\subseteq\mathbb{R}^{n}$. In this case, we can reformulate (weak-CE) as
@@
where Constraints (4c) and (4d) ensure that an optimal solution $x$ is feasible and lies in $\mathcal{D}$, while Constraints (4b) ensure that the solution $x$ is optimal for the objective vector $c$. Note that the problem has one constraint for every feasible solution $y\in\mathcal{X}\cap\{y:\hat{a}^{\top}y\geq\hat{b}\}$, i.e., the formulation can be of exponential size. Hence, we iteratively generate the constraints by finding the most violating constraint via a separation problem in each iteration. This procedure is the same as in [KB23] applied to our problem setup. The whole procedure can be found in Algorithm 1. The correctness follows from the same argumentation as in [KB23].
@@
$v,\hat{a},\hat{b},\hat{c},\mathcal{D},\mathcal{X},\mathcal{H}_{a,b}$, known upper bound $d^{*}$ for (weak-CE)
@@
Let $(a^{*},b^{*})\in\mathcal{H}_{a,b}$ be an optimal weak CE, i.e., an optimal solution of Problem (weak-CE). Then there exists a solution $x^{*}\in\mathcal{D}$ which is optimal for the problem
@@
The latter two lemmas indicate that we can solve Problem (weak-CE) by iterating over all values $v\in\mathcal{C}$ and solve the corresponding Problem (8). The best solution $(a^{*}(v),b^{*}(v))$ found over all these values must then be an optimal solution of (weak-CE). While this idea is feasible, it may happen that $\mathcal{C}$ contains a large number of values, and solving (weak-CE) for each is too expensive. To cut off unpromising solutions, we derive a lower bound for the objective function values in the following lemma.
Lemma 4.
@@
In this section, all parameters in the objective and the constraint are mutable. In this case the weak CE problem (weak-CE) can be formulated as follows:
@@
Since the number and type of constraints depend on the decision variables $a$ and $b$, the algorithms from the previous sections are not applicable. However, (weak-CE) can be reformulated as
@@ 4.2 Strong Counterfactual Explanations
In this subsection, only the objective parameters are mutable, i.e., $\hat{a}$ and $\hat{b}$ are fixed. The mutable parameter space is denoted as $\mathcal{H}_{c}\subset\mathbb{R}^{n}$. In this case, we can reformulate (strong-CE) as
@@
The same result of Lemma 3 holds for the strong CE version, i.e., there exists a $v^{*}\in\mathcal{C}$, such that any optimal solution of (12) is an optimal solution of (strong-CE). Likewise, similar to the lower bound in Lemma 4, it is easy to see that the following problem provides a lower bound for every problem (12) with integer $v\geq\bar{v}$:
@@
In this section all parameters in the objective and the constraints are mutable. In this case, the strong CE problem (strong-CE) can be reformulated as
@@ 5.1 Setup
We tested the algorithms on knapsack benchmark instances from kplib [^1] See https://github.com/likr/kplib., which we interpreted as cover instances to be consistent with the notation throughout this work. The instances were originally designed according to the procedure outlined in Kellerer, Pferschy & Pisinger [KPP04]. We used instances with a data range of 10001000222Note that Algorithms 3, 4, 6 and 7 are sensitive to changes in the data range by design, i.e., changes in the data range will lead to a proportional change in the number of subproblems, and thus the runtime. Since this is clear, we do not need to assess this empirically. and instance sizes in $[10,20,30,40]$, which were generated by slicing larger instances. Since it is well-known that the practical hardness of knapsack problems strongly depends on the instance structure [Pis05, SMCM21], e.g., if items value/weight ratios are strongly-correlated, the problems tend to be harder to solve in practice [Pis05], we use both $20$ uncorrelated and strongly correlated instances each. In total, we consider $4\times 2\times 20=160$ testing instances per favored solution space and CE type.
@@
Let $\mathcal{I}^{+},\mathcal{I}^{-}\subset\{1,\ldots,n\}$ with $\mathcal{I}^{+}\cap\mathcal{I}^{-}=\emptyset$ be index sets of variables that are to be fixed to $1$ or 0, respectively. Then, we define two sets for positive and negative fixations as
$\mathcal{D}^{+}=\{x\in\{0,1\}^{n}:x_{i}=1\ \forall i\in\mathcal{I}^{+}\}\text{ and }\ \mathcal{D}^{-}=\{x\in\{0,1\}^{n}:x_{i}=0\ \forall i\in\mathcal{I}^{-}\},$
respectively. In our computational experiments, the number of elements to be fixed is set to the following fraction:
@@
We also define a favored solution space $\mathcal{D}^{\geq}$ for given parameters $\alpha\in\mathbb{R}^{n}$ and $\beta\in\mathbb{R}$:
@@
We implemented all proposed algorithms in Python using Gurobi $12.0.0$ as a solver. All data and code are made available upon request. All computations were ran on RWTH Aachen Universities’ High-Performance Computing (HPC) CLAIX on Intel Xeon 8468 Sapphire nodes with a single $2.1$ GHz core and 55GB RAM each. We used a maximum runtime of 10h.
5.2 Results
@@
Although strongly correlated instances are generally considered to be harder, the fact that all items have a similar cost/weight ratio means small changes in weights are sufficient to add/remove an element from a solution. This is illustrated in the final subsection of the computational study. For feasible instances regardless of correlation, formulating a CE required on average a relative change of $0.25\pm 0.48\%$ in weights, with a range of between $0\%-3\%$. However, note that for many uncorrelated instances, we did not find a feasible solution. Thus, for those instances, a larger mutable parameter space might be necessary.
Note that a larger parameter spaces imply larger ranges of $\mathcal{C}$ as well. In our study, up to $2590$ different potential objective values were considered per instance with an average of $1160\pm 597$ values. This excludes the effect of the lower bound proposed in Lemma 4 shown in the next subsection.
@@
Figure 2: Progress towards optimal solution for $480$ strong CE instances of size $10,20,30$ that were solved to optimality. Red are primal and blue are dual bounds. The $x$-axis is normalized with respect to the total number of values considered in $\mathcal{C}$. The $y$-axis is normalized around each instance’s optimum.
Figure 3: Progress towards solution for $480$ weak CE instances of size $10,20,30$ that were solved to optimality. Red are primal and blue are dual bounds. The $x$-axis is normalized with respect to the total number of values considered in $\mathcal{C}$. The $y$-axis is normalized around each instance’s optimum.
We also observe that good solutions are centered around the halfway point of $50\%$, which tends to be close to the original objective value. Thus heuristically searching for small changes to the weights that lead to solutions close to the original objective function value appears to be a promising search strategy. However, this sacrifices the lower bound from Lemma 4 that significantly accelerates the exact approach.
Finally, we note that all the observations above apply for both weak and strong CEs equally.
@@
As to be expected, runtime increases significantly with instance size. Figures 4 and 5 illustrate this. The number of cuts required scales equally with instance size, as can be seen in Figures 6 and 7. In summary, CE can be computed using the algorithms provided. However, runtime and memory demands scale exponentially with instance size, which is to be expected for a $\Sigma_{2}^{p}$-complete problem.
Figure 4: Boxplot comparison of runtimes for $480$ strong CE instances with sizes $10$, 2020, $30$, and $40$. Instances not solved to optimality in time are assigned $36000$s/1010h.
Figure 5: Boxplot comparison of runtimes for $480$ weak CE instances with sizes $10$, 2020, $30$, and $40$. Instances not solved to optimality in time are assigned $36000$s/1010h.
Figure 6: Boxplot comparison of total number of generated cuts for $480$ strong CE instances with sizes $10$, 2020, $30$, and $40$.
Figure 7: Boxplot comparison of total number of generated cuts for $480$ weak CE instances with sizes $10$, 2020, $30$, and $40$.
@@
Computing CEs can also enable us to quantify (partial) decision values in terms of their contribution to an objective. As an example, consider CEs for the cover instance in Table 1.
@@
We observe multiple different types of modifications: To enforce item 0, the incumbent solution is made less attractive, especially item $7$, which is then substituted by items 0 and $6$. For items $4,5$ and $6$, we see that CEs themselves contain alternatives, e.g., to ensure either item is part of an optimal solution multiple equivalent changes of weights lead to the new solution of $[2,4,5,6]$. Item $8$ is significantly less attractive, enforcing it to be part of a solution requires a higher CE cost, even more so for items $3$ and $9$. Finally, we observe that simple metrics such as cost/weight are insufficient to capture modeling dynamics – item $9$ is very good in terms of cost/weight ratio, but still unattractive for optimal solutions. Thus, CEs allow for more differentiated evaluation of (partial) solutions.
@@ 6 Conclusion
In this work, we studied weak and strong counterfactual explanations for integer (linear) optimization problems. We showed that in the general case, computing a counterfactual increases complexity by one level in the polynomial hierarchy, i.e., $\mathsf{NP}$}-complete problems have a counterfactual that is $\Sigma_{2}^{p}$-complete, i.e., equivalent to general bilevel integer linear programming. We present algorithms to calculate counterfactual explanations for integer linear problems and implemented a tractable version thereof, which we evaluated on cover problem instances. We found that the proposed approach works well for instances with up to $40$ items.
@@
Acknowledgments
Computations were performed with computing resources granted by RWTH Aachen University High Performance computing. We thank Yosuke Onoue, who build and hosts the klib knapsack instance library. We thank Christina Büsing as well as Hector Geffner for enabling this research collaboration. One of the authors of this work, Felix Engelhardt, was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 2236/2.
@@ Appendix A Algorithms for Strong CEs
$v,\hat{a},\hat{b},\hat{c},\mathcal{D},\mathcal{X},\mathcal{H}_{a,b}$, known upper bound $d^{*}$ for (weak-CE)

### v2

@@ 4.1 Weak Counterfactual Explanations
If $\hat{a}$ and $\hat{b}$ are fixed, the mutable parameter space is denoted by $\mathcal{H}_{c}\subseteq\mathbb{R}^{n}$. In this case, we can reformulate (3) as
@@
where Constraints (7c) and (7d) ensure that an optimal solution $x$ is feasible and lies in $\mathcal{D}$, while Constraints (7b) ensure that the solution $x$ is optimal for the objective vector $c$. Note that the problem has one constraint for every feasible solution $y\in\mathcal{X}\cap\{y:\hat{a}^{\top}y\geq\hat{b}\}$, i.e., the formulation can be of exponential size. Hence, we iteratively generate the constraints by finding the most violating constraint via a separation problem in each iteration. This procedure is the same as in [korikov2023objective] applied to our problem setup. The whole procedure can be found in Algorithm 1. The correctness follows from the same argumentation as in [korikov2023objective].
@@
$v,\hat{a},\hat{b},\hat{c},\mathcal{D},\mathcal{X},\mathcal{H}_{a,b}$, known upper bound $d^{*}$ for (3)
@@
Let $(a^{*},b^{*})\in\mathcal{H}_{a,b}$ be an optimal weak CE, i.e., an optimal solution of Problem (3). Then there exists a solution $x^{*}\in\mathcal{D}$ which is optimal for the problem
@@
The latter two lemmas indicate that we can solve Problem (3) by iterating over all values $v\in\mathcal{C}$ and solve the corresponding Problem (11). The best solution $(a^{*}(v),b^{*}(v))$ found over all these values must then be an optimal solution of (3). While this idea is feasible, it may happen that $\mathcal{C}$ contains a large number of values, and solving (3) for each is too expensive. To cut off unpromising solutions, we derive a lower bound for the objective function values in the following lemma.
Lemma 4
@@
In this section, all parameters in the objective and the constraint are mutable. In this case the weak CE problem (3) can be formulated as follows:
@@
Since the number and type of constraints depend on the decision variables $a$ and $b$, the algorithms from the previous sections are not applicable. However, (3) can be reformulated as
@@ 4.2 Strong Counterfactual Explanations
In this subsection, only the objective parameters are mutable, i.e., $\hat{a}$ and $\hat{b}$ are fixed. The mutable parameter space is denoted as $\mathcal{H}_{c}\subset\mathbb{R}^{n}$. In this case, we can reformulate (5) as
@@
The same result of Lemma 3 holds for the strong CE version, i.e., there exists a $v^{*}\in\mathcal{C}$, such that any optimal solution of (15) is an optimal solution of (5). Likewise, similar to the lower bound in Lemma 4, it is easy to see that the following problem provides a lower bound for every problem (15) with integer $v\geq\bar{v}$:
@@
In this section all parameters in the objective and the constraints are mutable. In this case, the strong CE problem (5) can be reformulated as
@@ 5.1 Setup of Knapsack Instances
We tested the algorithms on knapsack benchmark instances from kplib [^1] See https://github.com/likr/kplib., which we interpreted as cover instances to be consistent with the notation throughout this work. The instances were originally designed according to the procedure outlined in Kellerer, Pferschy & Pisinger [kellerer2004exact]. We used instances with a data range of 10001000222Note that Algorithms 3, 4, 6 and 7 are sensitive to changes in the data range by design, i.e., changes in the data range will lead to a proportional change in the number of subproblems, and thus the runtime. and instance sizes in $[10,20,30,40]$, which were generated by slicing larger instances. Since it is well-known that the practical hardness of knapsack problems depends on the instance structure [Pisinger2005, SmithMiles2021], e.g., if items value/weight ratios are correlated, the problems tend to be harder to solve in practice [Pisinger2005], we use both $20$ uncorrelated and strongly correlated instances each. In total, we consider $4\times 2\times 20=160$ testing instances per favored solution space and CE type.
@@
Let $\mathcal{I}^{+},\mathcal{I}^{-}\subset\{1,\ldots,n\}$ with $\mathcal{I}^{+}\cap\mathcal{I}^{-}=\emptyset$ be index sets of variables that are to be fixed to $1$ or 0, respectively. Then, we define two sets for positive and negative fixations as $\mathcal{D}^{+}=\{x\in\{0,1\}^{n}:x_{i}=1\ \forall i\in\mathcal{I}^{+}\}$ and $\mathcal{D}^{-}=\{x\in\{0,1\}^{n}:x_{i}=0\ \forall i\in\mathcal{I}^{-}\}$ respectively. In our computational experiments, the number of elements to be fixed is set to the following fraction:
@@
We also define a favored solution space $\mathcal{D}^{\geq}$ for given parameters $\alpha\in\{0,1\}^{n}$ and $\beta\in\mathbb{N}$:
@@
We implemented all proposed algorithms in Python using Gurobi $12.0.0$ as a solver. All data and code are made available upon request. All computations were ran on RWTH Aachen Universities’ High-Performance Computing (HPC) CLAIX on Intel Xeon 8468 Sapphire nodes with a single $2.1$ GHz core and 55GB RAM each. We used a maximum runtime of 10h. A full implementation of all code is provided under [codetopaper]  [^3] See https://zenodo.org/records/18110317.
5.2 Computational Results on Knapsack Instances
@@
Finding a feasible solution required on average a relative change of $0.25\pm 0.48\%$ in weights, with a range of $0\%-3\%$. However, note that for many uncorrelated instances, we did not find a feasible solution. Thus, for those instances, a larger mutable parameter space might be necessary. Note that larger parameter spaces imply larger ranges of $\mathcal{C}$ as well. In our study, up to $2590$ different potential objective values were considered per instance with an average of $1160\pm 597$ values. This excludes the effect of the lower bound proposed in Lemma 4 shown in the next subsection.
@@
Figure 2: Progress towards optimal solution for $480$ strong CE instances of size $10,20,30,40$ that were solved to optimality. Red are primal and blue are dual bounds. The $x$-axis is normalized to the total number of values in $\mathcal{C}$. The $y$-axis is normalized around each instance’s optimum.
Figure 3: Progress towards solution for $480$ weak CE instances of size $10,20,30,40$ that were solved to optimality. Red are primal and blue are dual bounds. The $x$-axis is normalized to the total number of values in $\mathcal{C}$. The $y$-axis is normalized around each instance’s optimum.
We also observe that good solutions are centered around the halfway point of $50\%$, which tends to be close to the original objective value. Thus heuristically searching for small changes to the weights that lead to solutions close to the original objective function value appears to be a promising search strategy. However, this sacrifices the lower bound from Lemma 4 that significantly accelerates the exact approach. The observations above apply for both weak and strong CEs equally.
@@
As to be expected, runtime increases significantly with instance size. Figures 4(a) and 4(b) illustrate this. The number of cuts required scales equally with instance size, as can be seen in Figures 4(c) and 4(d). In summary, CE can be computed using the algorithms provided. However, runtime and memory demands scale exponentially with instance size, which is to be expected for a $\Sigma_{2}^{p}$-complete problem.
(a) Strong CE runtimes
(b) Weak CE runtimes
(c) Strong CE number of cuts
(d) Weak CE number of cuts
Figure 4: Boxplot comparison of runtimes (top row) and total number of generated cuts (bottom row) for $480$ strong and weak CE instances with sizes $10$, 2020, $30$, and $40$. Instances not solved to optimality are assigned $36000$s (1010h).
We can also illustrate this using primal-dual gaps. For that we restrict the data displayed to the 1st and 2nd item from each size, i.e $10$, 2020, $30$, and $40$. This is shown in Figure 5.
Figure 5: Progress towards solution for $64/96$ strong CE instances solved to optimality. Blue instances are size $40$, red instances are size $30$, yellow and green correspond to sizes $20,10$. The $x$-axis is normalized to the total number of values in $\mathcal{C}$. The $y$-axis is normalized around each instance’s optimum.
@@
Computing CEs can also enable us to quantify (partial) decision values in terms of their contribution to an objective. Although strongly correlated instances are generally considered to be harder, the fact that all items have a similar cost/weight ratio means small changes in weights are sufficient to add/remove an element from a solution.
As an example, consider CEs for the cover instance in Table 1.
@@
We observe different types of modifications: To enforce item 0, the incumbent solution is made less attractive, thus item $7$ is then substituted by items 0 and $6$. For items $4,5$ and $6$, we see that CEs themselves contain alternatives, e.g., to ensure either item is part of an optimal solution multiple equivalent changes of weights lead to the new solution of $[2,4,5,6]$. Item $8$ is significantly less attractive, enforcing it to be part of a solution requires a higher CE cost, even more so for items $3$ and $9$. Finally, we observe that simple metrics such as cost/weight are insufficient to capture modeling dynamics – item $9$ is very good in terms of cost/weight ratio, but still unattractive for optimal solutions. Thus, CEs allow for more differentiated evaluation of (partial) solutions.
5.3 An Application to the Resource Constrained Shortest Path Problem
In many cases, even considering few mutable constraints allows us to derive insights into an instance. Next, we examine the following setting: The task is to route traffic through a given graph, where using an arc comes with a toll. Given a limited toll budget this problem can be modeled as a Resource Constrained Shortest Path (RSCP). A CE in the toll parameters asks now how we have to change the toll such that an optimal path would use a certain edge or not.
Figure 6: CE for RCSP problem. Edge values indicate (length, toll). The red path is the original shortest path for a toll budget of 88; the least change in toll that leads to $e_{(1,2)}$ not being part of an optimal solution is given by lowering the toll of $e_{(3,4)}$ by one unit, resulting in the green RCSP.
Figure 6 illustrates a counterfactual in such a setting for a toy example. If we assume a toll budget of $8$ units, initially, the red path is the shortest path, with total length $12$. Assume the arc $e_{(1,2)}$ goes through the city center and we would like to avoid routing traffic through there whenever possible. Our question would then be: What is the smallest change so that the arc $e_{(1,2)}$ is no longer part of a shortest path? The result is given by the green path, which has total length $9$ and requires a single change to the data, i.e., lowering the toll of arc $e_{(3,5)}$ by one unit. In our example, this would mean that if we wanted to avoid routing traffic through the city center $e_{(1,2)}$, the best way to do so is to waive the toll on $e_{(3,5)}$ making it more attractive to bypass the city center.
In terms of mathematics, note that the RCSP contains multiple flow conservation constraints. However, since those are immutable, they do not affect our algorithm. Furthermore, we can bound the cost values $c$ by finding the optimal RCSP for each feasible value of $b$, i.e., the toll budget.
Solving RCSP pricing problems.
Finally, we test our algorithm on a set of benchmark instances from literature. For that, we use the SPPRCLIB data set provided by Jepsen et al. [jepsen2008branch]. The instances represent pricing problems from solving capacitated vehicle routing problems with column generation, see [Fukasawa2006]. As such, they contain negative node weights, as well as a resource constraint on the nodes. In Appendix B, we provide a ILP formulation of such problems. We restrict ourselves to the case that source and sink are the first and last indexed node, respectively, and round node weights to the nearest thousand and scaled down accordingly, e.g. $43000$ to $43$. The instances are complete graphs with 4526245-262 nodes. We investigate what would be the least change to the weights in the resource constraint by a maximum of $10\%$ to exclude the current first node from the shortest path. In the context of a pricing problem, this would imply computing a least change to the duals that would have led to a different column being added.
Of $44$ instances, we found a weak CE for $22$. Of these $22$ instances, for $5$ the original solution was already a counterfactual explanation. Finding a CE took an average of $15\pm 19$s and a maximum of $63$s. In about half the cases, the new RSCPs only have small alterations, for the other half, a completely new path is chosen. We note that we can thus compute CEs also for larger and more complex problems, provided the mutable part is limited. In the context of RSCP, this might allow insights into reoccurring patterns or the lack thereof in column generation.
@@ 6 Conclusion
In this work, we studied weak and strong counterfactual explanations (CEs) for integer (linear) optimization problems. We showed that in the general case, computing a counterfactual increases complexity by one level in the polynomial hierarchy, that is, $\mathsf{NP}$}-complete problems have a counterfactual that is $\Sigma_{2}^{p}$-complete, which is equivalent to general bilevel integer linear programming. We present algorithms to calculate counterfactual explanations for integer linear problems and implemented a tractable version thereof, which we evaluated on cover problem instances. We found that the proposed approach works well for instances with up to $40$ items.
@@
7 Code and Data Disclosure
The code and data to support the numerical experiments in this paper can be found at [codetopaper] , respectively under the link https://zenodo.org/records/18110317.
Acknowledgements
Computations were performed with computing resources granted by RWTH Aachen University High Performance computing. We thank Yosuke Onoue, who build and hosts the klib knapsack instance library. We thank Christina Büsing and Hector Geffner for enabling this research collaboration, as well as Christoph Grüne for insightful discussions on complexity theory.
@@ Appendix A Algorithms for Strong CEs
$v,\hat{a},\hat{b},\hat{c},\mathcal{D},\mathcal{X},\mathcal{H}_{a,b}$, known upper bound $d^{*}$ for (3)
@@
Appendix B Integer Programming Formulations for the Resource Constrained Shortest Path Problems
For the toy example, let $c_{ij}$ be costs for arc $ij\in A$ and $w_{ij}$ the respective weights. Furthermore, we designate a start node $s$ and a target node $t$. Then, we define a decision variable $x_{ij}\in\{0,1\}$ that is $1$ if arc $ij$ is part of the shortest path and 0 otherwise. The RCSPP is then given by:
min $\sum_{(i,j)\in A}c_{ij}x_{ij}$
s. t. $\sum_{j:(i,j)\in A}x_{ij}-\sum_{j:(j,i)\in A}x_{ji}=\begin{cases}1,&i=s,\\ -1,&i=t,\\ 0,&\text{else,}\end{cases}$ $\forall i\in V,$
$\sum_{(i,j)\in A}w_{ij}\,x_{ij}\leq b,$ (R)
$x_{ij}\in\{0,1\},$ $\forall(i,j)\in A.$
In our setting, we assume that only the resource constraint $R$ is mutable.
For the SPPRCLIB instances, the resource constraints are formulated on nodes, not edges. Furthermore, nodes also incur cost $c_{k}$ for a node $k$. Thus, we use a modified formulation, as seen below:
min $\sum_{(i,j)\in A}c_{ij}x_{ij}+\sum_{(i,j)\in A}\frac{c_{i}+c_{j}}{2}x_{ij}+\frac{c_{s}+c_{t}}{2}$
s. t. $\sum_{j:(i,j)\in A}x_{ij}-\sum_{j:(j,i)\in A}x_{ji}=\begin{cases}1,&i=s,\\ -1,&i=t,\\ 0,&\text{else,}\end{cases}$ $\forall i\in V,$
$\sum_{(i,j)\in A}\frac{w_{i}+c_{j}}{2}x_{ij}+\frac{w_{s}+w_{t}}{2}\leq b,$ (R)
$x_{ij}\in\{0,1\},$ $\forall(i,j)\in A.$